# Globally-Constrained Decentralized Optimization with Variable Coupling

### v3

@@ IV-C Parameter Selections
    Upper bound of $\|\bm{\lambda}^{\star}\|$: By virtue of the Slater’s condition of problem (6) and the continuity of $\sum_{i\in\mathcal{V}}f_{i}(x_{\mathcal{N}_{i}})$ over the compact $X_{\mathcal{V}}$, the method in [43] can be employed to derive an upper bound on $\|\bm{\lambda}^{\star}\|$.
@@
    where $\bar{A}=\operatorname{diag}(\bar{A}_{1},\cdots,\bar{A}_{n})$ with each $\bar{A}_{i}=(\sum_{j\in\mathcal{N}_{i}}A_{ji})^{T}$ and $\bar{\mathbf{u}}_{1}^{\star}=[(u_{11}^{\star})^{T},\cdots,(u_{n1}^{\star})^{T}]^{T}$. Since $W\mathbf{u}^{\star}=\mathbf{u}^{\star}$, we have $\|\mathbf{u}^{\star}\|^{2}_{W}=\|\mathbf{u}^{\star}\|^{2}=\sum_{i=1}^{n}(\|u_{i1}^{\star}\|^{2}+\|u_{i2}^{\star}\|^{2})=\|\bar{\mathbf{u}}_{1}^{\star}\|^{2}+\|\bm{\lambda}^{\star}\|^{2}$. Hence, it remains to bound $\|\bar{\mathbf{u}}_{1}^{\star}\|$. This may be realized by viewing (33) as a linear equation with $\bar{\mathbf{u}}_{1}^{\star}$ being the unknown, which is guaranteed to have a solution, and deriving the expression of any solution. For example, when each $\bar{A}_{i}$ has full column rank, which occurs in many real-world problems such as task offloading over mobile edge computing networks [26] and economic dispatch for power systems [44], the linear equation (33) has a unique solution
@@ V-A First Numerical Example
Problem (34) arises in, for example, the quality service of wireless networks and the plug-in electric vehicles charging problem [22, 23]. In this simulation, we consider a undirected network with $n=50$ nodes (see Figure 1), and the set of nodes is $\mathcal{V}=\{1,\ldots,50\}$. For each $i\in\mathcal{V}$, the constants $c_{i},d_{i}$ are chosen randomly from a uniform distribution in $[0,1]$ and $b=5$. Clearly, the linear objective and smooth constraints in problem (34) meet Assumption 1(a)–(e). The Slater’s condition can be satisfied when $x_{i}=1~{}\forall i\in\mathcal{V}$. Problem (34) satisfies not only Assumption 1 but also the assumptions for the dual subgradient method [20], $\text{DSA}_{2}$ [22] and C-SP-SG [23].
(a) Optimal error at $\mathbf{x}^{k}$
(b) Optimal error at $\bar{\mathbf{x}}^{k}$
Figure 2: Convergence performance of Algorithm 1 and three alternative methods for solving problem (34).
Based on above settings, we compare Algorithm 1 with the dual subgradient method [20], $\text{DSA}_{2}$ [22] and C-SP-SG [23] since they all have convergence rate guarantees for solving problem (34). For a fair comparison, we hand-tune all the algorithm parameters to make the algorithms achieve the best possible convergence performance. The optimal solution of problem (34) $\mathbf{x}^{\star}=[(x_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T}]^{T}$ is obtained by using the solver fmincon in Matlab. Figure 2 plots the optimal errors $\|\mathbf{x}^{k}-\mathbf{x}^{\star}\|$ and $\|\bar{\mathbf{x}}^{k}-\mathbf{x}^{\star}\|$, respectively, where $\mathbf{x}^{k}=[(x_{1}^{k})^{T},(x_{2}^{k})^{T},\ldots,(x_{n}^{k})^{T}]^{T}$ is the network-wide primal iterate at iteration $k$, and $\bar{\mathbf{x}}^{k}:=\frac{1}{k}\sum\limits_{\ell=1}^{k}\mathbf{x}^{\ell}$ is its running average. From Figure 2, we observe that Algorithm 1 converges faster than the other three algorithms, especially for the optimal error $\|\mathbf{x}^{k}-\mathbf{x}^{\star}\|$ at the primal iterate, demonstrating the superior convergence performance of our proposed algorithm.
@@ V-B Second Numerical Example
where each $P_{i},A_{i}\in\mathbb{R}^{d|\mathcal{N}_{i}|\times d|\mathcal{N}_{i}|}$ is symmetric positive semidefinite for guaranteeing the convexity of local functions. Besides, $Q_{i},a_{i}\in\mathbb{R}^{d|\mathcal{N}_{i}|}$, $B_{i}\in\mathbb{R}^{m\times d|\mathcal{N}_{i}|},b_{i}\in\mathbb{R}^{m}$, and $d_{i}\in\mathbb{R}$. These problem data are randomly generated under the above conditions. Under the same network environment as Figure 1, let $m=d=2$ and $X_{i}=\{x\in\mathbb{R}^{d}:l_{i}\leq x\leq u_{i}\}$, $f_{i}\left(x_{\mathcal{N}_{i}}\right)=(x_{\mathcal{N}_{i}})^{T}P_{i}x_{\mathcal{N}_{i}}+Q_{i}^{T}x_{\mathcal{N}_{i}}$ and $g_{i}\left(x_{\mathcal{N}_{i}}\right)=(x_{\mathcal{N}_{i}})^{T}A_{i}x_{\mathcal{N}_{i}}+a_{i}^{T}x_{\mathcal{N}_{i}}-d_{i}$. Apparently, problem (35) satisfies Assumption 1 under proper problem data.
@@
Problem (35) has a complicated coupling structure in the form of (1) and no existing decentralized optimization algorithms can solve (35). Hence, we only execute Algorithm 1 to solve it. Figure 3 depicts the optimal errors at $\mathbf{x}^{k}$ and $\bar{\mathbf{x}}^{k}$, respectively. The results in Figure 3 verify the effectiveness of Algorithm 1 on solving problems with both variable coupling and globally-coupled constraints.
@@ VI Conclusion
where the second step comes from $-2\langle g_{i}(x_{\mathcal{N}_{i}})-g_{i}(x_{\mathcal{N}_{i}}^{\prime}),t_{i}-t_{i}^{\prime}\rangle\leq\frac{1}{\beta^{2}}\|g_{i}(x_{\mathcal{N}_{i}})-g_{i}(x_{\mathcal{N}_{i}}^{\prime})\|^{2}+\beta^{2}\|t_{i}-t_{i}^{\prime}\|^{2}$ $\forall i\in\mathcal{V}$. Therefore, property (c) holds.
To prove property (d), let $\mathbf{x}^{\star}=[(x_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T}]$ be an optimum of problem (1), which is guaranteed to exist due to Assumption 1(f). Then, we let $t_{i}^{\star}=g_{i}(x_{\mathcal{N}_{i}}^{\star})-\frac{1}{n}\sum_{j\in\mathcal{V}}g_{j}(x_{\mathcal{N}_{j}}^{\star})$ $\forall i\in\mathcal{V}$, and let $\mathbf{y}^{\star}=[(x_{1}^{\star})^{T},(t_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T},(t_{n}^{\star})^{T}]^{T}$. It can be shown that $\mathbf{y}^{\star}$ is feasible to problem (6). Moreover, because $\mathbf{f}(\mathbf{y}^{\star})=\sum_{i\in\mathcal{V}}f_{i}(x_{\mathcal{N}_{i}}^{\star})$ and because (1) and (6) share the same optimal value, $\mathbf{y}^{\star}$ is an optimum of problem (6).
Finally, let $\tilde{x}_{i}$ $\forall i\in\mathcal{V}$ be given by Assumption 1(g), and let $\tilde{t}_{i}=g_{i}(\tilde{x}_{\mathcal{N}_{i}})-\frac{1}{n}\sum_{j\in\mathcal{V}}g_{j}(\tilde{x}_{\mathcal{N}_{j}})$ $\forall i\in\mathcal{V}$. It can thus be verified that $[(\tilde{x}_{1})^{T},(\tilde{t}_{1})^{T},\ldots,(\tilde{x}_{n})^{T},(\tilde{t}_{n})^{T}]^{T}$ is a Slater’s point for problem (6), so that property (e) holds.
@@
where $q_{i\ell}^{k}$ and $G_{i\ell}$, $\ell=1,\ldots,p$ represent the $\ell$th coordinates of $\mathbf{q}^{k}$ and $G_{i}$, respectively. Note from $\mathbf{q}^{0}=\max\{-\mathbf{G}(\mathbf{y}^{0}),\mathbf{0}\}$ and (8) that $\mathbf{q}^{k}+\mathbf{G}(\mathbf{y}^{k})\geq\mathbf{0}$. It follows from Assumption 1(d) that $\|(q_{i\ell}^{k}+G_{i\ell}(\mathbf{y}^{k}))(\nabla G_{i\ell}(\mathbf{y})-\nabla G_{i\ell}(\mathbf{y^{\prime}}))\|=(q_{i\ell}^{k}+G_{i\ell}(\mathbf{y}^{k}))\|\nabla G_{i\ell}(\mathbf{y})-\nabla G_{i\ell}(\mathbf{y^{\prime}})\|\leq(q_{i\ell}^{k}+G_{i\ell}(\mathbf{y}^{k}))L_{g}\|\mathbf{y}-\mathbf{y^{\prime}}\|$. Incorporating this and Proposition 1(b) into (39) leads to $\|\nabla\mathcal{R}^{k}(\mathbf{y})-\nabla\mathcal{R}^{k}(\mathbf{y^{\prime}})\|\leq L_{\mathcal{R}^{k}}\|\mathbf{y}-\mathbf{y^{\prime}}\|$. ∎
@@
It follows from [36, Lemma 3 and Lemma 4]. ∎
@@
To prove (42), we first bound $\mathbf{f}(\mathbf{y}^{k+1})-\mathbf{f}(\mathbf{y}^{\star})$ $\forall k\geq 0$. Recall from Section III that $(\mathbf{d}^{k})^{T}$ is the gradient of $\mathcal{R}^{k}(\mathbf{y})$ at $\mathbf{y}^{k}$. Thus, (7) can be rewritten as $\mathbf{y}^{k+1}=\operatorname{arg\;min}_{\mathbf{y}\in Y}M^{k}(\mathbf{y})$, where $M^{k}(\mathbf{y})=\mathcal{R}^{k}(\mathbf{y}^{k})+\nabla\mathcal{R}^{k}(\mathbf{y}^{k})^{T}(\mathbf{y}-\mathbf{y}^{k})+\frac{1}{2\gamma}\|\mathbf{y}-\mathbf{y}^{k}\|^{2}$. Since $M^{k}(\mathbf{y})$ is $\frac{1}{\gamma}$-strongly convex and $\mathbf{y}^{k+1}$ minimizes $M^{k}(\mathbf{y})$, we have $M^{k}(\mathbf{y}^{\star})\geq M^{k}(\mathbf{y}^{k+1})+\frac{1}{2\gamma}\|\mathbf{y}^{k+1}-\mathbf{y}^{\star}\|^{2}$, which is equivalent to
@@
By referring to the proof of [25, Lemma 2], we are able to derive the bounds on $M_{1}^{k}(\mathbf{y}^{\star})-M_{1}^{k}(\mathbf{y}^{k+1})$ and $M_{2}^{k}(\mathbf{y}^{\star})-M_{2}^{k}(\mathbf{y}^{k+1})$ in (45) as follows:
@@
By referring to the proof of [25, eq.(40), eq.(41)], we can show that $\sum_{k=1}^{K}\mathbf{G}(\mathbf{y}^{k})\leq\mathbf{q}^{K}$ and $\sum_{k=1}^{K}(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\mathbf{y}^{k}-\mathbf{c})=\rho(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{u}^{K}-\mathbf{u}^{0})$. Also note that $\bm{\lambda}^{\star}\geq\mathbf{0}$ and $\mathbf{u}^{\star}\in\operatorname{Range}(\mathbf{1}_{n}\otimes I_{m+p})$. It follows from (49) that
@@
In addition, following the proof of [25, Theorem 1], we can show via Proposition 1(c) that $\|\mathbf{G}(\mathbf{y}^{K})\|^{2}\leq\frac{1}{2}\|\mathbf{q}^{K}\|^{2}+\tilde{\beta}^{2}\|\mathbf{y}^{K}-\mathbf{y}^{\star}\|^{2}+\|\mathbf{G}(\mathbf{y}^{\star})\|^{2}$. By incorporating this into the expression of $S^{K}$ given in Lemma 4 and because $\frac{1}{\gamma}\geq\tilde{\beta}^{2}+\frac{1}{\rho}\|\mathbf{B}^{T}\mathbf{B}\|$,
@@
Incorporating (54) and (55) into (53) yields $\mathbf{f}(\bar{\mathbf{y}}^{k})-\mathbf{f}(\mathbf{y}^{\star})\geq-\frac{C^{0}}{k}$. This completes the proof.

### v4

@@ IV-C Parameter Selections
    Upper bound of $\|\bm{\lambda}^{\star}\|$: By virtue of the Slater’s condition of problem (6) and the continuity of $\sum_{i\in\mathcal{V}}f_{i}(x_{\mathcal{N}_{i}})$ over the compact $X_{\mathcal{V}}$, the method in [nedic2009approximate] can be employed to derive an upper bound on $\|\bm{\lambda}^{\star}\|$.
@@
    where $\bar{A}=\operatorname{diag}(\bar{A}_{1},\cdots,\bar{A}_{n})$ with each $\bar{A}_{i}=(\sum_{j\in\mathcal{N}_{i}}A_{ji})^{T}$ and $\bar{\mathbf{u}}_{1}^{\star}=[(u_{11}^{\star})^{T},\cdots,(u_{n1}^{\star})^{T}]^{T}$. Since $W\mathbf{u}^{\star}=\mathbf{u}^{\star}$, we have $\|\mathbf{u}^{\star}\|^{2}_{W}=\|\mathbf{u}^{\star}\|^{2}=\sum_{i=1}^{n}(\|u_{i1}^{\star}\|^{2}+\|u_{i2}^{\star}\|^{2})=\|\bar{\mathbf{u}}_{1}^{\star}\|^{2}+\|\bm{\lambda}^{\star}\|^{2}$. Hence, it remains to bound $\|\bar{\mathbf{u}}_{1}^{\star}\|$. This may be realized by viewing (33) as a linear equation with $\bar{\mathbf{u}}_{1}^{\star}$ being the unknown, which is guaranteed to have a solution, and deriving the expression of any solution. For example, when each $\bar{A}_{i}$ has full column rank, which occurs in many real-world problems such as task offloading over mobile edge computing networks [xiao2018distributed] and economic dispatch for power systems [yang2016distributed], the linear equation (33) has a unique solution
@@ V-A First Numerical Example
Figure 2: Convergence performance of Algorithm 1 and alternative methods for solving problem (34).
Problem (34) arises in, for example, the quality service of wireless networks and the plug-in electric vehicles charging problem [LiuC2020, Mateos2016]. In this simulation, we consider a undirected network with $n=50$ nodes (see Fig. 1), and the set of nodes is $\mathcal{V}=\{1,\ldots,50\}$. For each $i\in\mathcal{V}$, the constants $c_{i},d_{i}$ are chosen randomly from a uniform distribution in $[0,1]$ and $b=5$. Clearly, the linear objective and smooth constraints in problem (34) meet Assumption 1(a)–(e). The Slater’s condition can be satisfied when $x_{i}=1~\forall i\in\mathcal{V}$. Problem (34) satisfies not only Assumption 1 but also the assumptions in [LiangS2019, LiangS2019a, Mateos2016, falsone2023augmented, LiuC2020, wu2022distributed].
Under the above settings, we compare Algorithm 1 with the algorithms proposed in [LiangS2019, LiangS2019a, Mateos2016, falsone2023augmented, LiuC2020, wu2022distributed]. For a fair comparison, we hand-tune all the algorithm parameters to achieve the best possible convergence performance. The optimal solution of problem (34) $\mathbf{x}^{\star}=[(x_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T}]^{T}$ is obtained by using the solver fmincon in Matlab. Fig. 2 plots the sum of the objective error and constraint violation at $\mathbf{x}^{k}$ and $\bar{\mathbf{x}}^{k}$, respectively, i.e., $|\sum_{i=1}^{50}c_{i}x_{i}^{k}-c_{i}x_{i}^{\star}|+\|\max\{\sum_{i=1}^{50}-d_{i}\operatorname{log}(1+x_{i}^{k})+b,0\}\|$ and $|\sum_{i=1}^{50}c_{i}\bar{x}_{i}^{k}\!-\!c_{i}x_{i}^{\star}|+\|\max\{\sum_{i=1}^{50}-d_{i}\operatorname{log}(1+\bar{x}_{i}^{k})+b,0\}\|$, where $\mathbf{x}^{k}=[(x_{1}^{k})^{T},(x_{2}^{k})^{T},\ldots,(x_{n}^{k})^{T}]^{T}$ is the network-wide primal iterate at iteration $k$, and $\bar{\mathbf{x}}^{k}:=\sum\limits_{\ell=1}^{k}\mathbf{x}^{\ell}/k$ is its running average.
From Fig. 2, we observe that Algorithm 1 converges faster than the algorithms in [falsone2023augmented, LiangS2019a, LiuC2020, Mateos2016]. It also achieves a comparable or slightly better convergence rate compared to the algorithms in [LiangS2019] and [wu2022distributed]. Nevertheless, [LiangS2019] needs to transmit four times the number of variables as Algorithm 1 per iteration, leading to much higher communication cost, and [wu2022distributed] requires higher computational costs per iteration than Algorithm 1. The numerical results demonstrate the competitive convergence performance of our proposed algorithm.
@@ V-B Second Numerical Example
where each $P_{i},A_{i}\in\mathbb{R}^{d|\mathcal{N}_{i}|\times d|\mathcal{N}_{i}|}$ is symmetric positive semidefinite for guaranteeing the convexity of local functions. Besides, $Q_{i},a_{i}\in\mathbb{R}^{d|\mathcal{N}_{i}|}$, $B_{i}\in\mathbb{R}^{m\times d|\mathcal{N}_{i}|},b_{i}\in\mathbb{R}^{m}$, and $d_{i}\in\mathbb{R}$. These problem data are randomly generated under the above conditions. Under the same network environment as Fig. 1, let $m=d=2$ and $X_{i}=\{x\in\mathbb{R}^{d}:l_{i}\leq x\leq u_{i}\}$, $f_{i}\left(x_{\mathcal{N}_{i}}\right)=(x_{\mathcal{N}_{i}})^{T}P_{i}x_{\mathcal{N}_{i}}+Q_{i}^{T}x_{\mathcal{N}_{i}}$ and $g_{i}\left(x_{\mathcal{N}_{i}}\right)=(x_{\mathcal{N}_{i}})^{T}A_{i}x_{\mathcal{N}_{i}}+a_{i}^{T}x_{\mathcal{N}_{i}}-d_{i}$. Apparently, problem (35) satisfies Assumption 1 under proper problem data.
@@
Problem (35) has a complicated coupling structure in the form of (1) and no existing decentralized optimization algorithms can solve (35). Hence, we only execute Algorithm 1 to solve it. Fig. 3 depicts the objective error and the sum of inequality constraint violation and equality constraint violation at $\mathbf{x}^{k}$ and $\bar{\mathbf{x}}^{k}$, respectively. The results in Fig. 3 verify the effectiveness of Algorithm 1 on solving problems with both variable coupling and globally-coupled constraints.
@@ VI Conclusion
where the second step comes from $2\langle g_{i}(x_{\mathcal{N}_{i}})-g_{i}(x_{\mathcal{N}_{i}}^{\prime}),t_{i}-t_{i}^{\prime}\rangle\leq\frac{1}{\beta^{2}}\|g_{i}(x_{\mathcal{N}_{i}})-g_{i}(x_{\mathcal{N}_{i}}^{\prime})\|^{2}+\beta^{2}\|t_{i}-t_{i}^{\prime}\|^{2}$ $\forall i\in\mathcal{V}$. Therefore, property (c) holds.
To prove property (d), let $\mathbf{x}^{\star}=[(x_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T}]^{T}$ be an optimum of problem (1), which is guaranteed to exist due to Assumption 1(f). Then, we let $t_{i}^{\star}=g_{i}(x_{\mathcal{N}_{i}}^{\star})-\frac{1}{n}\sum_{j\in\mathcal{V}}g_{j}(x_{\mathcal{N}_{j}}^{\star})$ $\forall i\in\mathcal{V}$, and let $\mathbf{y}^{\star}=[(x_{1}^{\star})^{T},(t_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T},(t_{n}^{\star})^{T}]^{T}$. It can be shown that $\mathbf{y}^{\star}$ is feasible to problem (6). Moreover, because $\mathbf{f}(\mathbf{y}^{\star})=\sum_{i\in\mathcal{V}}f_{i}(x_{\mathcal{N}_{i}}^{\star})$ and because (1) and (6) share the same optimal value, $\mathbf{y}^{\star}$ is an optimum of problem (6). On the other hand, let $\mathbf{y}^{\star}=[(x_{1}^{\star})^{T},(t_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T},(t_{n}^{\star})^{T}]^{T}$ be an optimum of problem (6). We can verify that $(x_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T}]^{T}$ is feasible to problem (1), where the coupled equality constraint in (1) is directly satisfied and the coupled inequality constraint $\sum_{i=1}^{n}g_{i}(x_{\mathcal{N}_{i}}^{\star})\leq\mathbf{0}_{p}$ is obtained by summing $\sum_{i=1}^{n}t_{i}^{\star}=\mathbf{0}_{p}$ and each $g_{i}(x_{\mathcal{N}_{i}}^{\star})-t_{i}^{\star}\leq\mathbf{0}_{p}$ in (6). Moreover, $\sum_{i\in\mathcal{V}}f_{i}(x_{\mathcal{N}_{i}}^{\star})=\mathbf{f}(\mathbf{y}^{\star})$. Therefore, $(x_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T}]^{T}$ is an optimum of problem (1) and property (d) holds.
Finally, let $\tilde{x}_{i}$ $\forall i\in\mathcal{V}$ be given by Assumption 1(g), and let $\tilde{t}_{i}=g_{i}(\tilde{x}_{\mathcal{N}_{i}})-\frac{1}{n}\sum_{j\in\mathcal{V}}g_{j}(\tilde{x}_{\mathcal{N}_{j}})$ $\forall i\in\mathcal{V}$. It can thus be verified that $(\tilde{x}_{1})^{T},(\tilde{t}_{1})^{T},\ldots,(\tilde{x}_{n})^{T},(\tilde{t}_{n})^{T}]^{T}$ is a Slater’s point for problem (6), so that property (e) holds.
@@
[…]
@@
It follows from [YuHao2017, Lemma 3 and Lemma 4]. ∎
@@
To prove (42), we first bound $\mathbf{f}(\mathbf{y}^{k+1})-\mathbf{f}(\mathbf{y}^{\star})$ $\forall k\geq 0$. Recall from Section III that $\mathbf{d}^{k}$ is the gradient of $\mathcal{R}^{k}(\mathbf{y})$ at $\mathbf{y}^{k}$. Thus, (7) can be rewritten as $\mathbf{y}^{k+1}=\operatorname{arg\;min}_{\mathbf{y}\in Y}M^{k}(\mathbf{y})$, where $M^{k}(\mathbf{y})=\mathcal{R}^{k}(\mathbf{y}^{k})+\nabla\mathcal{R}^{k}(\mathbf{y}^{k})^{T}(\mathbf{y}-\mathbf{y}^{k})+\frac{1}{2\gamma}\|\mathbf{y}-\mathbf{y}^{k}\|^{2}$. Since $M^{k}(\mathbf{y})$ is $\frac{1}{\gamma}$-strongly convex and $\mathbf{y}^{k+1}$ minimizes $M^{k}(\mathbf{y})$, we have $M^{k}(\mathbf{y}^{\star})\geq M^{k}(\mathbf{y}^{k+1})+\frac{1}{2\gamma}\|\mathbf{y}^{k+1}-\mathbf{y}^{\star}\|^{2}$, which is equivalent to
@@
By referring to the proof of [wu2022distributed, Lemma 2], we are able to derive the bounds on $M_{1}^{k}(\mathbf{y}^{\star})-M_{1}^{k}(\mathbf{y}^{k+1})$ and $M_{2}^{k}(\mathbf{y}^{\star})-M_{2}^{k}(\mathbf{y}^{k+1})$ in (45) as follows:
@@
By referring to the proof of [wu2022distributed, eq.(40), eq.(41)], we can show that $\sum_{k=1}^{K}\mathbf{G}(\mathbf{y}^{k})\leq\mathbf{q}^{K}$ and $\sum_{k=1}^{K}(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{B}\mathbf{y}^{k}-\mathbf{c})=\rho(\mathbf{1}_{n}\otimes I_{m+p})^{T}(\mathbf{u}^{K}-\mathbf{u}^{0})$. Also note that $\bm{\lambda}^{\star}\geq\mathbf{0}$ and $\mathbf{u}^{\star}\in\operatorname{Range}(\mathbf{1}_{n}\otimes I_{m+p})$. It follows from (49) that
@@
In addition, following the proof of [wu2022distributed, Theorem 1], we can show via Proposition 1(c) that $\|\mathbf{G}(\mathbf{y}^{K})\|^{2}\leq\frac{1}{2}\|\mathbf{q}^{K}\|^{2}+\tilde{\beta}^{2}\|\mathbf{y}^{K}-\mathbf{y}^{\star}\|^{2}+\|\mathbf{G}(\mathbf{y}^{\star})\|^{2}$. By incorporating this into the expression of $S^{K}$ given in Lemma 4 and because $\frac{1}{\gamma}\geq\tilde{\beta}^{2}+\frac{1}{\rho}\|\mathbf{B}^{T}\mathbf{B}\|$,
@@
Incorporating (54) and ($5$) into (53) yields $\mathbf{f}(\bar{\mathbf{y}}^{k})-\mathbf{f}(\mathbf{y}^{\star})\geq-\frac{C^{0}}{k}$. This completes the proof.